OpenStudy (anonymous):

find the sum of the 15 terms in the geometric sequence 4,12,36,108

OpenStudy (raden):

use formula : ____________________ sn = a1(r^n-1)/(r-1) ____________________ with : a1=1st term = 4 r=ratio=a2/a1=12/4=3 and just put n=15

OpenStudy (anonymous):

Geometric series have a common multiplier between each sequential number. notice how 12/4=36/12=108/36=3 Find the ratio and the summation is just your first term times your ratio raised the number of terms you have. \[S=a*r^n\] where a is your first term, r is your ratio and n is the number of terms you have.

OpenStudy (anonymous):

i was wrong it should be \[S=a*r^{n-1}\]

OpenStudy (raden):

dont forget, divided by (r-1)

OpenStudy (anonymous):

if you divide by r-1, you get a wrong answer. try the first term with your equation. 4*(4^(1-1))/(3-1)=4*1/2=2 The first term is clearly 4 as the sequence shows but your /(r-1) reduces that by half. That half carries through the sequence making all summations reduced by 1/2

OpenStudy (raden):

re-check that question... we find the sum of the 15 terms, not find the 15th term

OpenStudy (anonymous):

ah you're correct, my bad. i was mistaken.

OpenStudy (anonymous):

thank you so much

OpenStudy (raden):

welcome